OpenStudy (aakashsudhakar):

I'm solving a partial differential equation, and I'm getting caught up in the integration step for the Fourier series. Can someone assist? The basic integral is written below.

OpenStudy (aakashsudhakar):

\[\int\limits_{0}^{\pi} (\sin x) ^2 (\sin nx) dx\]

OpenStudy (aakashsudhakar):

If it matters, consider "n" an arbitrary constant. It matters when concerning the Fourier series, which isn't related to the integral itself.

OpenStudy (dan815):

hmm lemme check something.. when n is not 1, is not all 0 is it

OpenStudy (aakashsudhakar):

This assumes that "n" goes from 1 to infinity. My given tip is to treat it like any constant, but I'm still unsure what to do.

OpenStudy (dan815):